A man of mass 60 kg is standing on a boat of mass 140 kg,which is at rest in still water. The man is initially at 20 m from the shore. He starts walking on the boat for 4 seconds with a constant speed of 1.5m/s towards the shore . The final distance of the man from the shore is what???

Grade:11

1 Answers

Arun
6 years ago
Speed of man in boat=1.5 m/s
Time taken=4 sec
So, the distance covered on the boat= speed x time=1.5 x 4=6 m
Now, boat will displace in opposite direction for making the position of center of mass undisturbed.
Suppose this displacement is x and its speed is v.
Now, by conservation of momentum
(140+60) x v=60 x 1.5
v=0.45 m/s
So displacement of boat=0.45 x 4=1.8 m
So the total distance of the man from the shore=20-6 +1.8=15.8 m